OpenStudy (anonymous):

9.Given the minimum value of the quadratic function f(x)=2x^2+px+q is 1 when x=-3,find the value of p and of q. @ganeshie8 @iambatman @Jhannybean @mathmath333

OpenStudy (anonymous):

i got p=12 and q=37 not sure whether my answer is right

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = 4x + p\] and the minimum value is when derivative is zero so we can say \( 4x + p = 0 \) and we know that \(x=-3 \) so \( p = 12 \)

OpenStudy (anonymous):

\[\frac{ dy }{ dx }~?\]

OpenStudy (anonymous):

it's derivative of that function,if u have not studied it...remember this point,the equation \( ax^2 + bx + c \) when \(a >0\) has a minimum point \[(\frac{ -b }{ 2a } , \frac{ 4ac - b^2 }{ 4a })\]

OpenStudy (anonymous):

here \(\large \frac{ -b }{ 4 } = -3\) so \( b = 12 \)

OpenStudy (anonymous):

can we use complete square method? @PFEH.1999

OpenStudy (anonymous):

this way is easier and the point i said u comes from square method but if u want to use it it would get confusing...

OpenStudy (anonymous):

bcoz i haven't learn \[(\frac{ -b }{ 2a },\frac{ 4ac-b^2 }{ 4a })\]

OpenStudy (anonymous):

well then,use square method...you'll reach the same thing.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i got p=12 and q=37 but i'm not sure whether i'm right. @PFEH.1999

OpenStudy (anonymous):

let me check...

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

would u check ur work again?

OpenStudy (anonymous):

i got p=12 and q=37 @PFEH.1999

OpenStudy (anonymous):

ok write ur solution i got q=19

OpenStudy (anonymous):

\[f(x)=(2x+\frac{ p }{ 2 })^2-\frac{ p^2 }{ 4 }+q\] \[=2(x+\frac{ p }{ 4 })^2-\frac{ p^2 }{ 4 }+q\] \[=2(x+3)^2-36+q\] \[=2(x+3)^2+1\] \[\frac{ p }{ 4 }=3\] \[p=12\] \[-36+q=1\] \[q=37\]

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (mathmath333):

put \(p=12\) and find \(q\)

OpenStudy (anonymous):

i got q=19 @mathmath333

OpenStudy (mathmath333):

me too got \(19\) check it with your book answer

OpenStudy (anonymous):

the book answer written q=19 @mathmath333

OpenStudy (mathmath333):

\(\Huge \bf done!\)

OpenStudy (anonymous):

can u show me the full working @mathmath333 ?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ f(x)=2x^2+px+q\\~\\ \text{f(x)=1,when x=-3, so}\\~\\ 1=2(-3)^2+p(-3)+q\\~\\ 3p-q=17-----\color{red}{(1)}\\~\\ f(x)=2x^2+px+q\\~\\ y=2x^2+px+q\\~\\ \frac{ dy }{ dx } = 4x + p\\~\\ 0 = 4x + p\\~\\ 0 = 4(-3) + p\\~\\ p=12\\~\\ \text{put p=12 in (1) and get "q"} }\end{align}\)

OpenStudy (anonymous):

\(\bf\huge\color{#ff0000}{T}\color{#ff2000}{h}\color{#ff4000}{a}\color{#ff5f00}{n}\color{#ff7f00}{k}~\color{#ffaa00}{Y}\color{#ffd400}{o}\color{#bfff00}{u}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) @mathmath333

OpenStudy (anonymous):

and @PFEH.1999

OpenStudy (anonymous):

@MARC_ sry alot ;) my internet connection suddenly broke up and i couldn't answer any more.. @mathmath333 thank you for helping him ;)

OpenStudy (anonymous):

It's okay. @PFEH.1999 :)